OpenStudy (some.random.cool.kid):

Factor (4x − 1)2 − (3x − 5)2. (x − 6)(7x + 4) 12(x − 1)(x + 5) (x + 4)(7x − 6) (x − 6)(12x − 1) this one has selections to it makes it easier because now I can plug in some of the choices.

OpenStudy (some.random.cool.kid):

@zepdrix

zepdrix (zepdrix):

Oh hmm, this one is a little trickier. I guess we have to expand everything out first.\[\large\rm (4x − 1)^2 − (3x − 5)^2\]Remember how to FOIL and stuff?

OpenStudy (some.random.cool.kid):

yeah sure... i remember.

zepdrix (zepdrix):

I'll do the second one :)\[\large\rm (4x − 1)^2 − (9x^2-30x+25)\]How bout the other one? What do we get? :d

OpenStudy (some.random.cool.kid):

uh you lost me how did those numbers arrive?

zepdrix (zepdrix):

\[\large\rm (3x-5)^2=(3x-5)(3x-5)\]Do you remember how to FOIL? :o

OpenStudy (some.random.cool.kid):

first outer inner last... bu twhat goes first?

OpenStudy (some.random.cool.kid):

the 4x? or the imaginary number?

zepdrix (zepdrix):

imaginary number? whut? 0_o

zepdrix (zepdrix):

\[\rm (3x-5)(3x-5)\quad=\quad3x\cdot3x ~+~ -5\cdot3x ~+~ 3x\cdot(-5) ~+~ -5(-5)\]\[\rm (3x-5)(3x-5)\quad=\quad 9x^2 ~+~ -15x ~+~ -15x ~+~ 25\]

OpenStudy (some.random.cool.kid):

ok

OpenStudy (some.random.cool.kid):

is first one by any chance 12x?

zepdrix (zepdrix):

\[\rm (4x-1)(4x-1)\quad=\quad 4x(4x) ~+~ 4x(-1) ~+~ -1(4x) ~+~ -1(-1)\] I'm not sure what you're asking...

OpenStudy (some.random.cool.kid):

mmm ok